OpenStudy (gnrfan):

Calculate the average rate of change for the function f(x) = −x4 + 4x3 − 2x2 + x + 1, from x = 0 to x = 1.

OpenStudy (gnrfan):

0 1 2 7

OpenStudy (gnrfan):

@iGreen. answers right above

OpenStudy (anonymous):

\(f(x) = -x^4 + 4x^3 - 2x^2 + x + 1\) Okay, when x = 0, y is 1. And when x = 1, y is 3. (0, 1), (1, 3) Plug it into the slope formula: \(m = \dfrac{3-1}{1-0}\) What's 3 - 1?

OpenStudy (gnrfan):

2

OpenStudy (anonymous):

Yes, and what's 1 - 0?

OpenStudy (gnrfan):

still 1 right

OpenStudy (anonymous):

Yes, so we have: \(\dfrac{2}{1}\) And what's 2 / 1?

OpenStudy (gnrfan):

what do you mean?

OpenStudy (anonymous):

Divide 2 / 1

OpenStudy (gnrfan):

2

OpenStudy (anonymous):

Yes, so that's the average rate of change.

OpenStudy (gnrfan):

thank you, can you do more please. I have two more i dont know